# CAPability: A Comprehensive Visual Caption Benchmark for Evaluating Both Correctness and Thoroughness

### v3

@@ A.2 Explanation for One Represents All Strategy
For each dimension, we only annotate one element, though there may be more than one element existing for some dimensions. Therefore, our annotations do not cover the whole visual content. But for those who try to cover the whole visual content, it is actually pretty hard to achieve this, as we mentioned in Sec. 3.1, the objects can be divided into almost infinite granularity. We focus on keeping the randomness of element selection, thus covering the whole visual content in a statistical sense, based on the law of large numbers. Therefore, we can get the ability to evaluate the thoroughness of the generated captions by calculating the hit.
@@
A.3 Details of Data Balance and Final Distribution
@@
The purpose of data balance is to suppress the long-tail distribution, thus ensuring there are a certain number of samples of different difficulties in the benchmark. Fig. A1 shows two examples of the comparison of pre-annotated data distribution and final human-annotated data distribution. For object number and camera angle dimension, we first randomly sample nearly 75K samples and conduct model pre-annotation. The 75K samples consist of approximately 1/3 SA-1B sam , 1/3 COYO-100M coyo700m , and 1/3 our crawled from multiple public datasets and websites. For object number dimension, we select out all samples with counting from 2 to 16 same objects within an image. As shown in Fig. A1 (left upper), the counting follows the long-tail distribution, there are fewer images with more objects within an image. Therefore, we conduct data-balanced sampling. Specifically, we separately sample images for different counts, thus forcing the number of each counts more balanced after the human correcting and filtering process. For camera angle dimension, the dutch angle data is rare, therefore we keep all dutch angle data, and sample the same number of other three categories. After the human correcting and filtering, the number of these categories varies slightly. The situation in other dimensions is also similar.
@@
As we consider the counts, categories, etc. for each dimension to conduct the data balance and do not consider the data source (i.e., from SA-1B, COYO-700M or crawled by ourselves) during this process, the final data source distribution for dimensions of object category, object number, object color, scene, camera angle, OCR varies. For spatial relation, we directly choose 1/2 CompreCap comprecap and 1/2 SA-1B images as SA-1B is more likely to contain high-resolution images with complex object relationship scene. For style, we choose all realistic images from SA-1B, and crawl animated, special effect, and old-fashioned images by ourselves, all art-related images are from Wikipaintings wikipaintings . For character identification, we use all images from public dataset, i.e., Wukong-100M wukong100m rather than crawling to ensure the proper copyright. For dynamic object number, we directly use data from VSI-Bench vsibench . For action and event, we directly use videos from Dream-1K dream1k . We crawl all videos for camera movement dimension by ourselves as there are few camera movement data in existing datasets. We thank all public datasets and benchmarks, their excellent images, videos, and annotations provides much convenience for building our CAPability.
@@
A.4 Discussion about the designed metrics
@@ B.1 Implementation Details
Qwen2VL and Qwen2.5VL. We keep the default minimum and maximum image pixels in package qwen_vl_utils, which is $4*28*28$, and $16384*28*28$, respectively. We also keep default video settings, the fps is set to $2.0$, the maximum frames are $768$, the minimum video pixel is $128*28*28$, and the maximum video pixel is $768*28*28$.
@@ Methods
Referring $\text{Referring Ratio}=\frac{|S(\text{COR})\cup S(\text{INC})|}{|S(\text{ALL})|}.$ (A5)
Evaluation stability. To validate the stability and robustness of our GPT-4-Turbo based evaluation method, we take the inferred caption of Gemini-1.5-pro as the example, run our evaluation 5 times, and the result is shown in Fig. A7. We tag the fluctuation range, i.e., the difference between the maximum and minimum scores, besides the data point. Fig. A7 shows our strong stability, and our average range of precision and recall are 1.1% and 1.0%, respectively. This demonstrates the reliability and interpretability of our evaluation method, which matches annotated elements in the generated captions.
Moreover, we introduce two other models, Gemini-2.0-flash and Qwen2.5-Max, to replace the GPT-4-Turbo in our evaluation pipeline. Both of them the most popular and powerful SOTA language models. We re-run the evaluation process with new evaluation models with the same evaluation prompts. The average result of evaluations for closed-source models and open-source 72B models is shown in Fig. A8. Though the judgment criteria for each model may be slightly different, leading to only a slight difference in the scores, but the evaluation results show high consistency as they keep the same rank of these MLLMs as GPT-4-Turbo. This further demonstrates the high reliability, interpretability, and stability of our evaluation methods.

### v4

@@ A.2 Explanation for One Represents All Strategy
"One represents all" is designed for object selection for object-related dimensions, text for OCR dimension, and action. We further build other dimensions related to attributes and relationships of objects (object color, object number, spatial relation) based on the selected objects. By aggregating results over a large number of samples with random pairing, we achieve statistical coverage across a broad spectrum of relationships, granularities, and contexts. Therefore, our design ensures that both single-object properties and inter-object relationships are robustly evaluated at the benchmark level, while providing a practical and scalable way to balance annotation cost and comprehensive coverage.
We focus on keeping the randomness of element selection, thus covering the whole visual content in a statistical sense, based on the law of large numbers. Therefore, we can get the ability to evaluate the thoroughness of the generated captions by calculating the hit. The details about our efforts to ensure the randomness are as follows. 1) To minimize bias and keep randomness in the pre-annotation stage, we utilize multiple SOTA models (GPT-4o, Gemini-1.5-pro, Qwen-VL-Max) to generate candidate objects and take the union of their outputs. We then use code (Python’s random library) to randomly select one object from this union. 2) During manual annotation, as humans tend to select the most obvious objects in the image, annotators are not allowed to reselect or change the target object if the pre-annotation is incorrect, thus mitigating the bias; they are only permitted to judge correctness and filter out incorrect samples. This ensures that neither model nor human bias influences which element is selected for annotation, thus keeping randomness.
A.3 Details about Human Subjectivity Controlling
We acknowledge that differences in human annotator judgment can introduce subjectivity, particularly in interpreting synonyms or resolving ambiguous cases. To reduce this, annotators are required to flag any samples for which they have low confidence. All low-confidence samples are then independently annotated by two additional annotators, and the final decision is determined by majority vote among the three annotations. We also have a spot check and verification process. For each dimension, we ask other annotators (or ourselves) to randomly select 20% of the samples for verification. If the annotation accuracy falls below 97%, we hold a meeting to highlight the incorrect annotations, revise all annotations for that dimension, and repeat this process until the annotation accuracy meets our requirements.
@@ A.3 Details about Human Subjectivity Controlling
A.4 Details of Data Balance and Final Distribution
@@
The purpose of data balance is to suppress the long-tail distribution, thus ensuring there are a certain number of samples of different difficulties in the benchmark. Fig. A1 shows two examples of the comparison of pre-annotated data distribution and final human-annotated data distribution. For object number and camera angle dimension, we first randomly sample nearly 75K samples and conduct model pre-annotation. The 75K samples consist of approximately 1/3 SA-1B sam, 1/3 COYO-100M coyo700m, and 1/3 crawled from multiple public datasets and websites. For object number dimension, we select all samples with counting from 2 to 16 same objects within an image. As shown in Fig. A1 (left upper), the counting follows the long-tail distribution, there are fewer images with more objects within an image. Therefore, we conduct data-balanced sampling. Specifically, we separately sample images for different counts, thus forcing the number of each count to be more balanced after the human correcting and filtering process. For camera angle dimension, the dutch angle data is rare, therefore we keep all dutch angle data, and sample the same number of the other three categories. After the human correcting and filtering, the number of these categories varies slightly. The situation in other dimensions is also similar.
@@
As we consider the counts, categories, etc. For each dimension to conduct the data balance and not consider the data source (i.e., from SA-1B, COYO-700M, or crawled by ourselves) during this process, the final data source distribution for dimensions of object category, object number, object color, scene, camera angle, OCR varies. For data of object-related dimensions (object category, object number, object color, spatial relation), global-related dimensions (scene and style), and a small part of camera angle and OCR dimensions, we collect the base data in a hybrid way. Our approach involves first performing pre-annotation for these dimensions on a large pool $P_{all}$ of images (100K). We then filter out those samples that are not suitable for the corresponding dimension independently, donated as $P^{pre}_{i}$ (nearly 40K - 80K), where $i$ represents each dimension. After that, we conduct balanced sampling and manual annotation, resulting in approximately 1K samples per dimension, donated as $P^{final}_{i}$. For spatial relation, we directly choose 1/2 CompreCap comprecap and 1/2 SA-1B images as SA-1B is more likely to contain high-resolution images with complex object relationship scenes. For style, we choose all realistic images from SA-1B, and crawl animated, special effect, and old-fashioned images by ourselves, all art-related images are from Wikipaintings wikipaintings. For character identification, we use all images from the public dataset, i.e., Wukong-100M wukong100m rather than crawling to ensure proper copyright. For dynamic object number, we directly use data from VSI-Bench vsibench. For action and event, we directly use videos from Dream-1K dream1k. We crawl all videos for camera movement dimension by ourselves, as there is little camera movement data in existing datasets. While there may be some overlap among pre-annotated $P^{pre}_{i}$, the average of overlapping samples in the final $P^{final}_{i}$ is 3.95%, as shown in Tab. A1. We thank all public datasets and benchmarks, their excellent images, videos, and annotations provide much convenience for building our CAPability.
Table A1: The overlap among each dimension.
Obj.
Cate.
Obj.
Num.
Obj.
Color
Spa.
Rel.
Scene
Cam.
Ang.
OCR Style
Cha.
Iden.
(D) Obj. Num.
Cam.
Mov.
Act./Event
11.3% 10.8% 10.5% 2.5% 4.6% 4.6% 2.6% 0.5% 0% 0% 0% 0%
@@
A.5 Details about Annotators
We employed 24 outsourcing annotators to complete the annotation task, including 14 female annotators and 10 male annotators. Their ages range from 24 to 30 years old, and all are from mainland China. Each annotator holds at least an associate’s or bachelor’s degree, demonstrating their expertise level. In total, the annotators labeled 23,824 samples, averaging approximately 993 samples per person. Among these, 3,323 samples were annotated by three different annotators, and the final result was determined through voting.
We specifically instructed annotators on how to handle such cases during the annotation process for each dimension. For example, in the color dimension, if the main object is primarily one color with minor secondary colors (such as a white airplane with some painted markings), only the primary color should be annotated. If the object exhibits multiple primary colors, annotators were asked to label up to three colors if the object can be clearly described within this limit; otherwise, the sample was filtered out. As a result, each object in our dataset is annotated with at most three colors.
A.6 Ethical Impact
The geographic and cultural backgrounds of the data are diverse. For data obtained from other public datasets, the distribution of potential bias in our benchmark largely follows that of the original datasets, as we performed random sampling. For the data we collected ourselves, there is a certain degree of cultural bias towards mainland China and East Asian cultures, although mainstream Western cultures are also represented to some extent. However, coverage of cultures from other regions is relatively limited.
A.7 Discussion about the Designed Metrics
@@ B.1 Implementation Details
Qwen2VL and Qwen2.5VL. We keep the default minimum and maximum image pixels in package qwen_vl_utils, which is 428284*28*28, and 16384282816384*28*28, respectively. We also keep default video settings, the fps is set to $2.0$, the maximum frames are $768$, the minimum video pixel is 1282828128*28*28, and the maximum video pixel is 7682828768*28*28.
@@ Methods
Table A3: The average metric of image dimensions and video dimensions.
Models LLaVA-OV-72B Qwen2VL-72B InternVL2.5-78B Qwen2.5VL-72B GPT-4o (0806) Gemini-1.5-pro Gemini-2.0-flash
image precision 81.4 82.2 78.2 80.7 84.4 81.1 84.6
video precision 59.5 62.9 55.6 65.0 67.6 68.7 67.3
image hit 55.3 57.5 57.9 62.1 65.2 67.7 64.5
video hit 26.9 29.2 22.4 33.7 36.8 43.9 37.6
Table A4: The PPL results of each generated caption by different models.
PPL models LLaVA-OV-72B Qwen2VL-72B InternVL2.5-78B Qwen2.5VL-72B GPT-4o (0806) Gemini-1.5-pro Gemini-2.0-flash
Qwen3 PPL 3.56 4.60 5.13 5.01 8.64 8.45 8.16
LLaMA3.1 PPL 4.89 6.70 7.75 7.54 11.29 10.76 10.68
$\text{Referring Ratio}=\frac{|S(\text{COR})\cup S(\text{INC})|}{|S(\text{ALL})|}.$ (A5)
Metric analysis between different modalities. Tab. A3 shows the metric analysis between image dimensions and video dimensions. Across all models, performance on image dimensions is substantially higher than on video dimensions, for both precision and hit metrics. Even the best-performing models (GPT-4o, Gemini series) show nearly 20% or greater drop in precision/hit from image to video dimensions. For example, all models perform well on object category, scene, OCR, and style for both precision and hit, but all models cannot achieve a satisfactory level on all video dimensions for hit. This shows that time series modeling and multi-frame information integration are still the main challenges of current MLLMs. It is true that some dimensions may be inherently more difficult to represent in video than in pictures. This may be due to the following two reasons: 1) There is more redundant content with more visual tokens input into MLLMs, handling the long sequence is likely more difficult than a shorter one. 2) Temporal change and movement should be considered for video dimensions, which may lead to confusion for models to recognize. Nevertheless, images can also represent the strength of videos in this dimension to a certain extent (such as OCR), because video understanding also requires a good spatial understanding ability as a prerequisite. The video dimensions we designed are more focused on the challenge of temporality. In the future, we will consider introducing more video data to fully evaluate the video’s ability in the spatial dimensions.
@@
The naturalness and coherence analysis. In the current landscape of MLLMs, the naturalness and coherence of generated captions are generally no longer a sufficient or even primary differentiator of model performance. Linguistic quality is more considered for NLP language models, and modern LLMs consistently generate grammatically correct and fluent sentences chang2024survey. As MLLMs are built from LLMs, the main challenge and research focus of MLLMs has shifted toward evaluating the factual accuracy and relevance of the generated captions, rather than their linguistic quality. For example, the papers of MLLMs do not report any metrics about the naturalness and coherence, such as PPL llavaov; qwen2.5vl, modern benchmarks for MLLMs do not measure the naturalness and coherence either mmmu; videomme. To support the view that the naturalness and coherence of generated captions are not the primary challenges, we also further calculate perplexity (PPL) of the generated captions of each model. PPL is a fundamental metric for language models that gauges how “surprised” the model is by a given text, thus reflecting their naturalness and fluency, the lower PPL means better coherence. Specifically, we conduct qwen3-32B and LLaMA-3.1-8B-Instruct to calculate the PPL of generated captions from these models, as shown in Tab. A4. (There seems to be some gap between the closed-source APIs and open-source models. This may be because the training data distribution among open-source models is more similar to Qwen3 and LLaMA3.1 than closed-source APIs.) Among closed-source APIs, the PPLs of them are similar to or lower than GPT-4o. Among open-source models, the PPLs of them are similar to or lower than Qwen2.5VL-72B. Based on the common sense of GPT-4o and Qwen2.5VL-72B can generate coherent and human-like sentences, we can draw a conclusion that the naturalness and coherence are not the main challenge for all these models.
Evaluation stability. To validate the stability and robustness of our GPT-4-Turbo-based evaluation method, we take the inferred caption of Gemini-1.5-pro as the example, run our evaluation 5 times, and the result is shown in Fig. A7. We tag the fluctuation range, i.e., the difference between the maximum and minimum scores, besides the data point. Fig. A7 shows our strong stability, and our average range of precision and recall are 1.1% and 1.0%, respectively. This demonstrates the reliability and interpretability of our evaluation method, which matches annotated elements in the generated captions. Moreover, we introduce three other models, Gemini-2.0-flash, Qwen2.5-Max, and Claude3.7-Sonnet to replace the GPT-4-Turbo in our evaluation pipeline. All of them are the most popular and powerful SOTA language models. We re-run the evaluation process with new evaluation models with the same evaluation prompts. The average result of evaluations for closed-source models and open-source 72B models is shown in Fig. A8. Though the judgment criteria for each model may be slightly different, leading to only a slight difference in the scores, the evaluation results show high consistency as they keep the same rank of these MLLMs as GPT-4-Turbo. This further demonstrates the high reliability, interpretability, and stability of our evaluation methods.